# Explicit Multimodal Graph Modeling for Human-Object Interaction Detection

### v1

@@ Abstract
Transformer-based methods have recently become the prevailing approach for Human-Object Interaction (HOI) detection. However, the Transformer architecture does not explicitly model the relational structures inherent in HOI detection, which impedes the recognition of interactions. In contrast, Graph Neural Networks (GNNs) are inherently better suited for this task, as they explicitly model the relationships between human-object pairs. Therefore, in this paper, we propose Multimodal Graph Network Modeling (MGNM) that leverages GNN-based relational structures to enhance HOI detection. Specifically, we design a multimodal graph network framework that explicitly models the HOI task in a four-stage graph structure. Furthermore, we introduce a multi-level feature interaction mechanism within our graph network. This mechanism leverages multi-level vision and language features to enhance information propagation across human-object pairs. Consequently, our proposed MGNM achieves state-of-the-art performance on two widely used benchmarks: HICO-DET and V-COCO. Moreover, when integrated with a more advanced object detector, our method demonstrates a significant performance gain and maintains an effective balance between rare and non-rare classes.
Introduction
@@
Human-object interaction detection is a task that aims to localize human-object pairs and simultaneously infer the interactions between them. This task represents a more profound level of visual scene understanding than standard object detection. A detected HOI instance is represented as a triplet in the form of $\langle\text{human, action, object}\rangle$. Recently, HOI detection has garnered significant attention from the computer vision community, owing to its extensive applications in downstream tasks such as action recognition (Kong and Fu 2022), image caption (Hossain et al. 2018), and visual scene understanding (Ma et al. 2024a).
@@
Similar to object detectors, HOI detectors can be categorized into two primary types: one-stage and two-stage methods. Recent one-stage methods usually employ the Transformer architecture. Specifically, these methods often adapt DETR-like frameworks (Carion et al. 2020) to detect human-object pairs and predict the interactions in parallel via Transformer decoders (Liao et al. 2022; Zhou et al. 2022). However, compared with two-stage methods, one-stage methods cannot easily leverage the features from object detectors, potentially limiting their performance.
@@
Two-stage methods typically follow a distinct training paradigm, which generally involves fine-tuning a pre-trained object detector and subsequently freezing its parameters (Zhang, Campbell, and Gould 2022; Zhang et al. 2023). This paradigm allows the interaction predictor to focus solely on interaction prediction, thereby improving training efficiency. As illustrated in Figure 1(a), recent two-stage methods usually adopt Transformer to extract the visual features. However, the attention mechanism inherent in the Transformer updates features via a weighted-sum aggregation, which does not explicitly model the relational structure in HOI detection and thus hinders further performance gains. Motivated by this limitation, we propose a multimodal graph network framework designed to explicitly model the HOI detection task, which more closely aligns with the human perception of HOI.
@@
As illustrated in Figure 1(b), our GNN-based method begins by matching detected human and object instances to construct explicit relational connections for each human-object pair. Notably, as observed in prior studies (Zhang, Campbell, and Gould 2021; Park, Park, and Lee 2023), relying solely on the visual features of human and object nodes often fails to capture fine-grained contextual cues. Inspired by the success of Vision Language Models (VLMs) (Lei, Wang, and Tan 2024; Guo et al. 2024), we leverage CLIP (Radford et al. 2021) to generate rich multimodal features, thereby enhancing feature interaction between each human-object pair. Benefiting from the sufficient prior knowledge of CLIP, our method not only achieves a significant performance gain but also effectively mitigates the performance disparity between rare and non-rare classes, which is a persistent challenge for previous methods (Zhang, Campbell, and Gould 2021, 2022; Liao et al. 2020). Moreover, we introduce a Multi-level Feature Interaction (MFI) mechanism designed to further utilize these multimodal features and enhance the exchange of contextual information. This mechanism leverages both low-level and high-level vision and language features to intensify the interaction between human-object pairs, thereby enriching the graph network with more fine-grained semantic information.
@@
In light of the preceding analysis, we propose Multimodal Graph Network Modeling (MGNM), a novel and efficient two-stage framework for HOI detection. Our MGNM achieves state-of-the-art performance on two widely used benchmarks: HICO-DET (Chao et al. 2018) and V-COCO (Gupta and Malik 2015). Furthermore, when integrated with a stronger object detector, our method demonstrates a significant performance gain and maintains an effective balance between rare and non-rare classes. The effectiveness of each proposed component is validated through extensive ablation studies. The contributions of this work are as follows:
@@
  • We design a multi-level feature interaction mechanism that effectively fuses low-level and high-level semantic cues across modalities, substantially enhancing the exchange of fine-grained contextual information within the graph structure.
@@
  • Extensive experiments on HICO-DET and V-COCO benchmarks demonstrate that our method not only achieves state-of-the-art performance but also maintains an effective balance between rare and non-rare classes.
Related Work
One-stage Methods
@@
One-stage methods are characterized by their parallel detection of human and object instances and prediction of their corresponding interactions. Early approaches typically relied on heuristics such as human keypoints (Liao et al. 2020) or predefined union boxes (Kim et al. 2020). More recently, Transformer-based architectures have become the dominant paradigm for one-stage HOI detection. For instance, Chen et al. (Chen et al. 2021) employ two separate Transformer decoders to predict human and object instances and their interactions in parallel. Similarly, Yang et al. (Yang et al. 2023) leveraged VLMs to extract fine-grained semantic features, which were then processed by two sub-decoders to simultaneously predict HOI triplets. Nevertheless, one-stage methods are often encumbered by significant memory usage and slow convergence speed.
Two-stage Methods
@@
Two-stage methods first employ an off-the-shelf object detector to identify human and object instances, and subsequently predict interactions between the matched human-object pairs. Recent two-stage methods often use Transformer encoders to extract features or enhance attention mechanisms (Zhang, Campbell, and Gould 2022; Zhang et al. 2023). Another line of work (Guo et al. 2024; Lei et al. 2023; Lei, Wang, and Robby T. 2024) has incorporated VLMs for feature refinement, leading to notable accuracy improvements. However, while the Transformer excels as a powerful feature extractor, it does not offer an explicit relational modeling framework for the HOI detection task (Antoun and Asmar 2023). To address this gap, our method employs a multimodal graph network framework instead. Our work is situated within a growing line of GNN-based HOI detectors. For instance, SCG (Zhang, Campbell, and Gould 2021) utilized a GNN to facilitate information exchange between human and object nodes, while ViPLO (Park, Park, and Lee 2023) incorporated features from the CLIP image encoder and ViTPose (Xu et al. 2022) to enrich this interaction. Nevertheless, a key limitation of these methods is their primary reliance on visual features, which overlooks the rich semantic cues offered by multimodal features, thereby limiting their ultimate performance.
Method
@@
Figure 2: Overview of our MGNM framework.The proposed MGNM framework initiates by employing the DETR model to detect human and object instances, which are subsequently matched to form candidate human-object pairs. The features of these pairs are then processed through a four-stage multimodal graph network, consisting of the spatial, visual, textual, and interaction stages. (1) Spatial Stage: For each candidate pair, low-level spatial features derived from their bounding boxes are used to initialize pairwise representations. Simultaneously, these spatial features are utilized to construct a weighted adjacency matrix that regulates feature interactions in the subsequent stages. (2) Visual Stage: High-level visual semantic features are extracted using the CLIP image encoder. These features further enrich the interactions among human-object pairs by incorporating detailed visual context. (3) Textual Stage: The CLIP text encoder is adopted to obtain semantic embeddings for the corresponding human and object categories, providing complementary textual cues for each pair. (4) Interaction Stage: In the final stage, the model captures high-level interaction features between human-object pairs, facilitating more effective relational reasoning within the graph structure. Finally, the refined pairwise representations are utilized as queries in a Transformer decoder, which attends to global image features from the backbone (serving as keys and values), to predict the final HOI triplets.
@@
Following the standard two-stage paradigm for HOI detection, we assume that a given image contains $m$ detected humans and $n$ detected objects, resulting in $m\times n$ candidate human-object pairs. Here, we denote $\mathbf{N}\in\mathbb{R}^{(m+n)\times d}$ as the human and object instances, which serves as the node representations in the GNN. The corresponding human and object sets within the human-object pairs are represented as $\mathbf{H}\in\mathbb{R}^{(m\times n)\times d}$ and $\mathbf{O}\in\mathbb{R}^{(m\times n)\times d}$, respectively. $d$ is the dimension of the nodes. Additionally, the features of human-object pairs are denoted as $\mathbf{E}\in\mathbb{R}^{(m\times n)\times 2d}$, which can be interpreted as the edge embeddings within the GNN framework.
Multi-level Feature Interaction Mechanism
In this paper, we propose a novel Multi-level Feature Interaction (MFI) mechanism. This mechanism utilizes the low-level and high-level vision and language features to enhance the interactive modeling of human-object pairs. As illustrated in Figure 2, the MFI mechanism is structured into four sequential stages: Spatial, Visual, Textual, and Interaction. The details of each stage are elaborated below.
@@
The spatial stage is designed to exploit geometric relationships to enhance interactions between human-object pairs. Specifically, the spatial features comprise bounding box coordinates, box areas, Intersection over Union (IoU), and other related metrics (Zhang, Campbell, and Gould 2021). These features encode the fundamental geometric configurations of each human-object pair and serve as a source of low-level visual information. Moreover, since these are handcrafted features, their computation incurs minimal overhead for the model (Zhang, Campbell, and Gould 2022). Upon extracting the spatial features, we utilize a Multimodality Fusion (MMF) mechanism (Zhang et al. 2023) to integrate them with the initial human and object features, thereby generating the initial representation for each human-object pair.
$\mathbf{E}$ $=\text{MMF}(\mathbf{H}\oplus\mathbf{O},\mathbf{S}),$ (1)
where $\oplus$ means the concatenation operation and $\mathbf{S}$ means the spatial features. In addition, based on the spatial features and corresponding human and object features, the weight matrix $\mathbf{W}_{G}$ for the multimodal graph network is attained with the multi-branch fusion (MBF) mechanism (Zhang, Campbell, and Gould 2021).
$\mathbf{W}_{G}$ $=\text{Linear}(\text{MBF}(\mathbf{H}\oplus\mathbf{O},\mathbf{S})),$ (2)
where the linear layer is used to align the dimensions.
@@
The visual stage leverages the CLIP to incorporate high-level vision features. Specifically, we employ the CLIP image encoder to extract high-level visual semantic features from the given image. These visual features are then used to pass the high-level visual semantic information to the humans and objects as follows:
$\text{M}_{\mathcal{V}\rightarrow\mathcal{N}}(\mathbf{V},\mathbf{N})$ $=\sigma(\mathbf{W}_{G}\odot\text{MBF}_{v}(\mathbf{V},\mathbf{N})),$ (3)
where $\text{M}_{\mathcal{V}\rightarrow\mathcal{N}}$ denotes the message passing function responsible for propagating visual information to the human and object node features. $\mathbf{V}$ represents the visual features from CLIP. $\sigma$ is ReLU activation function and $\odot$ is element-wise multiplication. $\text{MBF}_{v}$ is a subclass of the MBF mechanism, tasked with integrating visual and node features. Finally, the human and object node features are updated using a layer normalization step, as formulated below:
$\mathbf{N}$ $=\text{LN}(\mathbf{N}+\text{M}_{\mathcal{V}\rightarrow\mathcal{N}}(\mathbf{V},\mathbf{N})),$ (4)
@@
The textual stage is dedicated to generating textual embeddings for each node, providing a source of low-level language features. As illustrated in Figure 2, we adopt the prompt template “a photo of a $\langle\text{object}\rangle$”, where the $\langle\text{object}\rangle$ comes from the detection results. In this paper, for brevity, we denote the humans and objects as the whole node set $\mathbf{N}$, and the corresponding textual features are assumed to be matched. The process is formulated as follows:
$\text{M}_{\mathcal{T}\rightarrow\mathcal{N}}(\mathbf{T},\mathbf{N})$ $=\sigma(\mathbf{W}_{G}\odot\text{MBF}_{t}(\mathbf{T},\mathbf{N})),$ (5)
where $\text{M}_{\mathcal{T}\rightarrow\mathcal{N}}$ means the message passing function responsible for propagating textual information to the human and object node features. $\mathbf{T}$ means the textual features from CLIP. $\text{MBF}_{t}$ is a subclass of the MBF mechanism. Similarly, we also use layer normalization step when updating the human and object features.
@@
To incorporate high-level interaction semantic cues, we again employ the CLIP text encoder. This time, it is used to generate interaction-centric semantic features, which constitute high-level language features. The prompt template “a photo of a person interacting with $\langle\text{object}\rangle$” is adopted, which aims to boost the feature interaction between human-object pairs. This process is formulated as follows:
$\mathbf{E}$ $=\text{LN}(\mathbf{E}+\mathbf{I}),$ (7)
@@
$\mathbf{E}$ $=\text{LN}(\mathbf{E}+\mathbf{H}\oplus\mathbf{O}).$ (8)
Multimodal Graph Network Modeling

### v2

@@ Abstract
Transformer-based methods have recently become the prevailing approach for Human-Object Interaction (HOI) detection. However, the Transformer architecture does not explicitly model the relational structures inherent in HOI detection, which impedes the recognition of interactions. In contrast, Graph Neural Networks (GNNs) are inherently better suited for this task, as they explicitly model the relationships between human-object pairs. Therefore, in this paper, we propose Multimodal Graph Network Modeling (MGNM) that leverages GNN-based relational structures to enhance HOI detection. Specifically, we design a multimodal graph network framework that explicitly models the HOI task in a four-stage graph structure. Furthermore, we introduce a multi-level feature interaction mechanism within our graph network. This mechanism leverages multi-level visual and language features to enhance information propagation across human-object pairs. Consequently, our proposed MGNM achieves state-of-the-art (SOTA) performance on two widely used benchmarks: HICO-DET and V-COCO. Moreover, when integrated with a more advanced object detector, our method demonstrates a significant performance gain and maintains an effective balance between rare and non-rare classes.
1 Introduction
@@
Human-object interaction detection is a task that aims to localize human-object pairs and simultaneously infer the interactions between them. This task represents a more profound level of visual scene understanding than standard object detection. A detected HOI instance is represented as a triplet in the form of $\langle\text{human, action, object}\rangle$. Recently, HOI detection has garnered significant attention from the computer vision community [wang2024interaction, Jia2024OrchestratingTS, Yang2025NoMS, Zhang2025HOIDR1RL, Xue2025GuidingHI, 10658050], owing to its extensive applications in downstream tasks such as action recognition [kong2022human, do2024skateformer], image caption [Hossain2018ACS, rotstein2024fusecap], and visual scene understanding [ma2024scene].
@@
Similar to object detectors, HOI detectors can be categorized into two primary types: one-stage and two-stage methods. Recent one-stage methods usually employ the Transformer architecture [Vaswani2017AttentionIA]. Specifically, these methods often adapt DETR [Carion2020EndtoEndOD] to detect human-object pairs and predict the interactions in parallel via Transformer decoders [Zhou2022HumanObjectID, Jia2024ContextHOISC]. However, compared with two-stage methods, one-stage methods cannot easily leverage the features from object detectors, potentially limiting their performance.
@@
Two-stage methods typically follow a distinct training paradigm, which generally involves fine-tuning a pre-trained object detector and subsequently freezing its parameters [zhang2022upt, zhang2023pvic, park2023viplo]. This paradigm allows the interaction predictor to focus solely on interaction prediction, thereby improving training efficiency. As illustrated in Figure 1(a), recent two-stage methods usually adopt Transformer to extract the visual features. However, the attention mechanism inherent in the Transformer updates features via a weighted-sum aggregation, which does not explicitly model the relational structure in HOI detection and thus hinders further performance gains. Motivated by this limitation, we propose a multimodal graph network framework designed to explicitly model the HOI detection task, which more closely aligns with the human perception of HOI.
@@
As illustrated in Figure 1(b), our GNN-based method begins by matching detected human and object instances to construct explicit relational connections for each human-object pair. Notably, as observed in prior studies [zhang2021scg, park2023viplo], relying solely on the visual features of human and object instances often fails to capture fine-grained contextual cues. Inspired by the success of Vision Language Models (VLMs) [Liao2022GENVLKTSA, Lei2024EZHOIVA, Guo2024UnseenNM, Ning2023HOICLIPEK], we leverage CLIP [Radford2021LearningTV] to generate rich multimodal features, thereby enhancing feature interaction between each human-object pair. Benefiting from the sufficient prior knowledge of CLIP, our method not only achieves a significant performance gain but also effectively mitigates the performance disparity between rare and non-rare classes, which is a persistent challenge for previous methods [zhang2021scg, zhang2022upt, liao2020ppdm, Ma2023FGAHOIFA]. Moreover, we introduce a Multi-level Feature Interaction (MFI) mechanism designed to further utilize these multimodal features and enhance the exchange of contextual information. Collaborating with the designed object- and interaction-centric prompts, this mechanism leverages both low-level and high-level visual and language features to intensify the feature interaction between human-object pairs, thereby enriching the graph network with more fine-grained semantic information.
@@
In light of the preceding analysis, we propose Multimodal Graph Network Modeling (MGNM), a novel and effective two-stage framework for HOI detection. Our MGNM achieves SOTA performance on two widely used benchmarks: HICO-DET [chao2018learning] and V-COCO [gupta2015visual]. Furthermore, when integrated with a more advanced object detector, our method demonstrates a significant performance gain and maintains an effective balance between rare and non-rare classes. The effectiveness of each proposed component is validated through extensive ablation studies. The contributions of this work are as follows:
@@
  • Collaborating with the object- and interaction-centric prompts, we design a multi-level feature interaction mechanism that effectively fuses low-level and high-level semantic cues across modalities, substantially enhancing the exchange of fine-grained contextual information within the graph structure.
@@
  • Extensive experiments on HICO-DET and V-COCO benchmarks demonstrate that our method not only achieves SOTA performance but also maintains an effective balance between rare and non-rare classes.
2 Related Work
2.1 One-stage Methods
@@
One-stage methods are characterized by their parallel detection of human and object instances and prediction of their corresponding interactions. Early approaches typically relied on heuristics such as human keypoints [liao2020ppdm, Zhong2021GlanceAG] or predefined union boxes [kim2020uniondet, Gkioxari2017DetectingAR]. More recently, Transformer-based architectures have become the dominant paradigm for one-stage HOI detection. For instance, Chen et al. [chen_2021_asnet] employ two separate Transformer decoders to predict human and object instances and their interactions in parallel. Similarly, Yang et al. [yang2023boosting] leveraged VLMs to extract fine-grained semantic features, which were then processed by two sub-decoders to simultaneously predict HOI triplets. Nevertheless, one-stage methods are often encumbered by significant memory usage and slow convergence speed.
2.2 Two-stage Methods
@@
Two-stage methods first employ an off-the-shelf object detector to identify human and object instances, and subsequently predict interactions between the matched human-object pairs. Recent two-stage methods often use Transformer to extract features or enhance attention mechanisms [zhang2022upt, zhang2023pvic]. Another line of work [Guo2024UnseenNM, 11092421, lei2024efficient, NEURIPS2024_2a54def4] has incorporated VLMs for feature refinement, and subsequently adopted Transformer decoders to fuse the features. However, while the Transformer excels as a powerful feature extractor, it does not offer an explicit relational modeling framework for the HOI detection task. To address this gap, our method employs a multimodal graph network framework instead. Our work is situated within a growing line of GNN-based HOI detectors [Qi2018LearningHI, Gao2020DRGDR, zhang2021scg, park2023viplo]. For instance, SCG [zhang2021scg] utilized a GNN to facilitate information exchange between human and object nodes, while ViPLO [park2023viplo] incorporated features from the CLIP image encoder and ViTPose-L [xu2022vitpose] to enrich this feature interaction. Nevertheless, a key limitation of these methods is they usually treat GNN as a feature extractor, but don’t analyze it from the perspective of explicit modeling. Moreover, they primarily rely on visual features and overlook the rich semantic cues offered by multimodal features, thereby limiting their further performance improvement.
3 Method
@@
Figure 2: Overview of our MGNM framework. The core of our proposed MGNM framework is a four-stage multimodal graph network. (1) Spatial Stage: For each candidate pair, low-level spatial features derived from their bounding boxes and 3D location prior are used to initialize pairwise representations. After that, these spatial features are utilized to construct a weighted matrix that regulates feature interactions in the subsequent stages. (2) Visual Stage: High-level visual semantic features are extracted using the CLIP image encoder. These visual cues further enrich the interactions among human-object pairs. (3) Textual Stage: The CLIP text encoder is adopted to obtain semantic cues for the corresponding human and object instances with the designed object-centric prompt. (4) Interaction Stage: In the final stage, collaborating with the interaction-centric prompt, the model captures high-level interaction features between human-object pairs, facilitating more effective relational reasoning within the graph structure. Finally, the refined pairwise representations are utilized as queries in the Transformer decoder to predict the final HOI triplets.
@@
In this paper, we assume a given image contains $n$ detected objects (including $m$ humans). With a standard one-to-one matching strategy, we obtain $m\times n$ candidate human-object pairs, which can be seen as the adjacency matrix in GNN. Here, we denote $\mathbf{N}\in\mathbb{R}^{n\times d}$ as the human and object instances set, which serves as the node representations in GNN. The corresponding human and object sets within the human-object pairs are represented as $\mathbf{H}\in\mathbb{R}^{(m\times n)\times d}$ and $\mathbf{O}\in\mathbb{R}^{(m\times n)\times d}$. $d$ is the dimension of the nodes. Additionally, the features of human-object pairs are denoted as $\mathbf{E}\in\mathbb{R}^{(m\times n)\times 2d}$, which can be interpreted as the edge embeddings within the GNN framework.
3.1 Multi-level Feature Interaction Mechanism
In this paper, we propose a novel Multi-level Feature Interaction (MFI) mechanism. This mechanism utilizes the low-level and high-level visual and language features to enhance the interactive modeling of human-object pairs. As illustrated in Figure 2, the MFI mechanism is structured into four sequential stages: Spatial, Visual, Textual, and Interaction. The details of each stage are elaborated below.
@@ 3.1 Multi-level Feature Interaction Mechanism
The spatial stage is designed to exploit geometric relationships to enhance interactions between human-object pairs. Specifically, the spatial features comprise bounding box coordinates, box areas, Intersection over Union (IoU), and other related metrics [zhang2021scg, zhang2022upt]. More details about spatial features can be found in the supplementary materials. These features encode the basic geometric relationships of each human-object pair and serve as a source of low-level visual information. Moreover, since these are handcrafted features, their computation incurs minimal overhead for the model [zhang2022upt]. Upon extracting the spatial features, we utilize a multimodality fusion (MMF) mechanism [zhang2023pvic] to integrate them with the initial human and object features, thereby generating the initial representations for each human-object pair:
$\mathbf{E}$ $=\text{MMF}\left(\mathbf{H}\oplus\mathbf{O},\mathbf{S}\right),$ (1)
where $\oplus$ means the concatenation operation and $\mathbf{S}$ means the spatial features. Then, inspired by [11092421], we add the 3D location prior ($\mathbf{L}$) into pair representations. As demonstrated by [11092421], these 3D location prior features are beneficial for the model to discriminate between with interaction and no interaction. The formulation is as follows:
$\mathbf{E}$ $=\text{MLP}\left(\mathbf{E}\oplus\mathbf{L}\right).$ (2)
In addition, based on the spatial features and corresponding human and object features, the weight matrix $\mathbf{W}_{G}\in\mathbb{R}^{m\times n}$ for the multimodal graph network is attained with the multi-branch fusion (MBF) mechanism [zhang2021scg]:
$\mathbf{W}_{G}$ $=\text{Linear}\left(\text{MBF}\left(\mathbf{H}\oplus\mathbf{O},\mathbf{S}\right)\right),$ (3)
where the linear layer is used to align the dimensions. More details about the MBF mechanism can be found in the supplementary materials.
@@
The visual stage leverages the CLIP to incorporate high-level visual features. Specifically, we employ the CLIP image encoder to extract high-level visual semantic features from the given image. These visual features are then used to pass the high-level visual semantic cues to humans and objects as follows:
$\text{M}_{\mathcal{V}\rightarrow\mathcal{N}}\left(\mathbf{V},\mathbf{N}\right)$ $=\text{ReLU}\left(\mathbf{W}_{G}\odot\text{MBF}_{v}\left(\mathbf{V},\mathbf{N}\right)\right),$ (4)
where $\text{M}_{\mathcal{V}\rightarrow\mathcal{N}}$ denotes the message passing function responsible for propagating visual information to the human and object node features. $\mathbf{V}$ represents the visual features from CLIP. $\odot$ is element-wise multiplication. $\text{MBF}_{v}$ is a subclass of the MBF mechanism, tasked with integrating visual and node features and providing complementary cues for the nodes. Finally, the human and object node features are updated using a layer normalization step, as formulated below:
$\mathbf{N}$ $=\text{LN}\left(\mathbf{N}+\text{M}_{\mathcal{V}\rightarrow\mathcal{N}}\left(\mathbf{V},\mathbf{N}\right)\right),$ (5)
@@
The textual stage is dedicated to generating textual embeddings for each node, providing a source of low-level language features. As illustrated in Figure 2, we adopt a simple object-centric prompt template “a $\langle\text{object}\rangle$ in the human-centered scene”, focusing on the expected object. The $\langle\text{object}\rangle$ comes from the detection results. In this paper, for brevity, we denote the humans and objects as the whole node set $\mathbf{N}$, and the corresponding textual features are assumed to be matched. The process is formulated as follows:
$\text{M}_{\mathcal{T}\rightarrow\mathcal{N}}(\mathbf{T},\mathbf{N})$ $=\text{ReLU}\left(\mathbf{W}_{G}\odot\text{MBF}_{t}\left(\mathbf{T},\mathbf{N}\right)\right),$ (6)
where $\text{M}_{\mathcal{T}\rightarrow\mathcal{N}}$ means the message passing function responsible for propagating textual information to the human and object node features. $\mathbf{T}$ means the textual features from CLIP. $\text{MBF}_{t}$ is a subclass of the MBF mechanism. Similarly, we also use the layer normalization step when updating the human and object features.
@@
To incorporate high-level interaction semantic cues, we propose the interaction-centric prompt. Specifically, this prompt is constructed using geometric and spatial cues closely related to the interaction, including IoU, relative size, relative position, and center distance. Based on the specific values of these cues, we dynamically generate distinct prompts for each human-object pair. For IoU, we define four descriptive phrases: “closely interacting with”, “directly engaging with”, “potentially interacting with”, and “near but may not be interacting with”. Similarly, for relative size, descriptive terms such as “very large”, “large”, and “small” are employed. By combining these dynamically selected phrases, the complete interaction-centric prompt is formed. For instance, an $\text{IoU}>0.3$ triggers the phrase “closely interacting with”, while a $\text{relative size}>3.0$ appends “very large $\langle\text{object}\rangle$”, where $\langle\text{object}\rangle$ is the predicted object class. The resulting prompt becomes: “a person closely interacting with very large $\langle\text{object}\rangle$”. In this situation, given the high IoU implying a direct interaction, explicit position and distance cues are deemed unnecessary. Conversely, $\text{IoU}\leq 0.05$, yields the phrase “near but may not be interacting with”. If this is combined with a $\text{relative size}\leq 0.3$, the prompt “very small $\langle\text{object}\rangle$” is added. Subsequently, a $\text{center distance}\in(0.15,0.3)$ appends positional descriptors such as “right” and “above”. Finally, the complete prompt becomes: “a person near but may not be interacting with very small $\langle\text{object}\rangle$ positioned above and to the right”. A comprehensive definition of these prompt-generation rules can be found in the supplementary materials. This interaction-centric prompt design provides discriminative semantic cues for the text encoder.
Leveraging this interaction-centric prompt, we once again employ the CLIP text encoder. This time, it is used to generate interaction-centric semantic features, which constitute high-level language features. This process is formulated as follows:
$\mathbf{E}$ $=\text{LN}\left(\mathbf{E}+\mathbf{I}\right),$ (8)
@@
$\mathbf{E}$ $=\text{LN}\left(\mathbf{E}+\mathbf{H}\oplus\mathbf{O}\right).$ (9)
3.2 Multimodal Graph Network Modeling